Solve this system of three equations in three unknowns:
1) x + y − z = 4
2) x − 2 y + 3 z = −6
3) 2 x + 3 y + z = 7

Answer by Boreal(15235)

add 1 and 3 to eliminate z
3x+4y=11
add 3 times 1 to 2 to also eliminate z
4x+y=6
now multiply the second by -4 and add to the top one
3x+4y=11
-16x-4y=-24
-13x=-13
x=1
therefore, y=2
and z=-1
(1, 2, -1)